### 09 Jan 2024 v2

# Token-free LLMs Can Generate Chinese Classical Poetry with More Accurate Format

Abstract.
Finetuned large language models (such as ChatGPT and Qwen-chat) can generate Chinese classical poetry following human’s instructions. LLMs perform well in content, but are usually lacking in format, with occasionally excess or insufficient number of characters in each line.
Since most SOTA LLMs are token-based, we assume that the format inaccuracy is due to the difficulty of the ”token planning” task, which means that the LLM need to know exactly how much characters are contained in each token and do length-control planning based on that knowledge.
In this paper, we first confirm our assumption by showing that existing token-based large language models has limited knowledge on token-character relationship. We use a spelling bee probing procedure, and find that Qwen-chat failed in nearly 15% Chinese spelling test.
We then show that a token-based model can be easily tailored into a token-free model (in terms of Chinese), which can largely solve the format accuracy problem. Our tailoring procedure removes long-tokens from the vocabulary and the language model head, and keeps only character-level or byte-level tokens.
As part of our contribution, we release the finetuned token-free model (which is based on Qwen-chat-7B), which can generate chinese classical poetry following complex instructions like LLMs (such as story paraphrasing), and also perform well in format. On the test set, our token-free model achives an format accuracy of 0.96, compared to 0.84 for token-based equivalents and 0.38 for GPT-4.
Token Free, Controllable Text Generation, Large Language Models, Classical Poetry
[…]
1. Introduction
Chinese classical poetry is a treasure in the Chinese language and Chinese culture. Finetuned large language models (such as ChatGPT(OpenAI, 2016) and Qwen-chat(Bai et al., 2023)) can now generate Chinese classical poetry following human’s instructions. These models perform well in content, but are usually lacking in format, with occasionally excess or insufficient number of characters in each line.
To further address on the problem, we first briefly introduce the format requirements of Chinese classical poetry. For detailed introduction, refer to (Hu and Sun, 2020).
Format requirements of Chinese classical poetry. Chinese classical poetry can be classified into two primary categories, SHI and CI, both have strict format requirements.
SHI: There are four kinds of SHI in total, their name and format requirements are listed below:
  1. (1)
    Wuyan Jueju (4 lines in total, 5 characters each line.)
  2. (2)
    Wuyan Lvshi (8 lines in total, 5 characters each line.)
  3. (3)
    Qiyan Jueju (4 lines in total, 7 characters each line.)
  4. (4)
    Qiyan Lvshi (8 lines in total, 7 characters each line.)
CI: CI is more complex, with nearly one thousand forms in total, each with different number of lines and specific requirement of number of characters for each line. One example is given below (This form is called Rumengling):
Requirement:
6 characters; 6 characters; 5 characters; 6 characters; 2 characters; 2 characters; 6 characters;
LLM Generated poetry. With the power of LLM, the instructions for poetry generation could be as simple as a provided keyword, or as complex as telling a story. One example of keyword-based generation with GPT-4 is given below.
In this example, GPT-4 is asked to write a poem the form of ”Rumengling”, with the key word ”cheerful”. We can see that the generated poem is relative in content, but clearly violate the format requirement, even with an exmaple format provided in the prompt. To present it clearer to the readers, the redundant characters are marked with a strikethrough.
Prompt 请写一首如梦令,关键词为“兴高采烈”。请严格按照如梦令对每一句话的字数要求,下面给出一个例子: 常记溪亭日暮,沉醉不知归路。 兴尽晚回舟,误入藕花深处。 争渡,争渡,惊起一滩鸥鹭。 Äpgjy Response 浅笑嫣然花开,兴高采烈醉春 半醺半醉半睡眠,悠然步入花丛 漫步,漫步,蝴蝶舞动春光
Prompt(translated into English) Please write a poem in the form of “Rumengling”, with the keyword “cheerful”. Please strictly follow the number of character requirements for each sentence. Here is an example: I often recall the sun setting at the riverside pavilion, lost in the intoxication and unaware of the way back. Later on when my excitement wanes, I return on the boat, only to find myself unwittingly entering lotus in clusters. Struggling to cross, struggling to cross, with Seagulls and herons startled by me and flew away.Äpgjy Response(translated into English) With a gentle smile, the flowers bloom, in high spirits, drunk on the spring breeze. Half drunk and half drowsy, leisurely I walk into the midst of the flowers. Strolling, strolling, with butterflies dancing in the thick spring light.
In fact, we tested 1000 prompts, with 100 keywords and 10 format styles for each keyword, and find that only 38% of the generated poems exactly met the format requirements.
Difficulty of the “token planning” task. Most SOTA LLMs are token-based: text is splited into tokens, with each token containing one or more bytes or characters. This means that the models have no direct access to the characters behind the tokens, and this is also the case in regard to Chinese.
In the light of this, we assume that the format inaccuracy is due to the difficulty of the ”token planning” task, which means that the LLM need to know exactly how much characters are contained in each token and do length-control planning based on that knowledge. The difficulty of ”token planning” task comes from two aspects:
  1. (1)
    Insufficient knowledge in token-character relationship. The models may be lacking in the exact knowledge of token-character relationship, so they are inaware of the exact number of characters that have been generated and to be generated.
  2. (2)
    Inability to sum and plan. Even if the model know the exact character composition of each token, it may make mistakes at summing and planing.
In the paper, we mainly focus on the first aspect. In section III, we will show that the LLM’s knowledge in token-character relationship is clearly limited, which is the key reason for the bad performance on format requirements.
Can token-free models do better? If token planning is the problem, can token-free models do better? Token-free model refer to the language models that tokenize text only in character-level or byte-level. Previous work has explored poem generation with token-free models, both in the English field (Belouadi and Eger, 2023) and the Chinese field (Hu and Sun, 2020). Those language models are pretrained from scratch.
In this paper, we use a different approach. Instead of training a token-free model from scratch, we start from existing LLMs to take advantage of their existing ability in instruction understanding and text generation. We tailor existing models in terms of vocabulary and language model head. We keep only character-level and byte-level tokens, transforming them into token-free models.
We use Qwen-chat-7B(Bai et al., 2023) as our base model, and tailor it into token-free model. The tailored model are then further finetuned with poetry instruction dataset. On the test set, our token-free model achives an format accuracy of 0.96, compared to 0.84 for token-based finetuned Qwen-chat-7B and 0.38 for GPT-4.
2. Related work
Figure 1. Tailor token-based model into token-free model.
Previous work(Kaushal and Mahowald, 2022; Itzhak and Levy, 2022) has addressed that the token-character relationship of LLMs may be limited. They designed and implemented probing procedure with the vocabulary embedding layer of language models to investigate their knowledge in token-character relationship. In regard to Chinese classical poetry generation with LLMs, further work should be done in two aspects:
  1. (1)
    The probing procedure should be done in Chinese.
  2. (2)
    All model parameters instead of only the vocabulary embedding layer should be considered. If we only consider the vocabulary embedding layer, the knowledge in later layers will be omitted and the knowledge level would be underestimated.
Existing general LLMs can generate Chinese poetry generation following instructions. There are also models designed specially for Chinese poetry generation task. We differ from those work in regard of methods and model ability. Detailed comparisons with the most related work is listed below.
2.1. Comparison with existing general LLMs
All SOTA general LLMs, including (OpenAI, 2016; Bai et al., 2023) are token-based. With our token-free architecture, format accuracy in poetry generation is greatly enhanced. We achives an format accuracy of 0.96 on our test set, compared to 0.84 for token-based counterpart and 0.38 for GPT-4.
2.2. Comparison with ByGPT5, Jiuge-GPT2 and other poetry generating models
Previous work has adopted token-free architecture for poetry generation including (Zhang and Lapata, 2014; Wang et al., 2016; Yan, 2016; Zhang et al., 2017; Yi et al., 2017, 2018; Zhipeng et al., 2019; Belouadi and Eger, 2023; Hu and Sun, 2020). Those models are pretrained from scratch, while our model are tailored from existing LLMs. As a result, our model inherits the ability to deal with more complex task from LLMs, such as generating poetry with main content (story paraphrasing).
3. Experiment 1: Probing into LLM’s knowledge in token-character relationship
In previous sections, we proposed an assumption that the format inaccuracy comes from the difficulty of the ”token planning” task, which means that the LLM need to know exactly how much characters are contained in each token and do length-control planning based on that knowledge. We will verify this assumption in this section.
Standard large language models use subword tokenization algorithms, like BPE(Sennrich et al., 2016) or WordPiece tokenization(Schuster and Nakajima, 2012), that partition text into word pieces. These tokenization algorithms face an apparent limitation in that they lack direct access to the character composition of each token. For tasks that demand a high degree of precision in character count, such as poetry writing, there is a necessity for precisely control over the length of generated content. The fulfillment of character count requirements can only be assured provided that the large model possesses an exceedingly accurate understanding of the character composition of tokens.
3.1. Method
We designs a procedure to probe large language models to thoroughly explore the character knowledge involved in them. Specifically, we probe whether the model has the knowledge that the token ”大模型” contains three characters ”大”, ”模” and ”型”. This probing procedure is called ”spelling bee” in previous work (Itzhak and Levy, 2022).
The models we investigate here are the Qwen-series (Qwen-7B-chat, Qwen-14B-chat and Qwen-72B-chat)(Bai et al., 2023). Qwen-series are the earliest open source LLM with strong ability in Chinese and are influencial in the Chinese community. The largest open source model in this series is 72B, which enables us to investigate how token-character knowledge evolves as model gets larger.
In the context of large language models, the probing procedure could be formulated as the following prompt template. Here the special sign $<|extra\_1|>$ is used to ensure long-token and single-character will both be tokenized as is.
Prompt List all the charaters in the following token: $<|extra\_1|>$大模型$<|extra\_1|>$pgjy Response $<|extra\_1|>$$<|extra\_1|>$$<|extra\_1|>$ $<|extra\_1|>$
In the context of large language models, we simply use supervised finetune (Ouyang et al., 2022) as the probing procedure. We randomly sample 1000 tokens from the vocabulary as test, and the remaining vocabulary are as training examples.
Similar to (Itzhak and Levy, 2022), in an effort to counteract the effects of similar tokens within the train and test data, we implement similarity filter, where we remove all related characters in the test set from the training set. For example, if the token ”大模型” is found in the test data, then any tokens containing the character ”大”, ”模” or ”型” would be eliminated from the train data.
Our procedure differs from previous work in two aspects:
  1. (1)
    Our experiments are designed to probe all language model parameters, while previous work(Kaushal and Mahowald, 2022; Itzhak and Levy, 2022) probes only the vocabulary embedding layer, which ignores the knowledge contained in later layers, and thus would underestimate the real knowledge level.
  2. (2)
    With the rapid advancement of large language models with a decoder-only architecture such as Qwen(Bai et al., 2023), we are able to conduct probing experiments relatively easily with the model itself through supervised finetune(Ouyang et al., 2022), while previous work probes with a new seperate model, which may be difficult to train and would also underestimate the real knowledge level.
To sum up, we belive that our procedure can better estimate how LLM’s knows about token-character relationship. At the same time, we replicate the same procedure as previous work (Itzhak and Levy, 2022), only as a reference.
3.2. Results
The results of the spelling bee probing procedures are summarized with failure rate, which is defined as the number of times when LLM fail to correctly output the character sequence of the required token, divided by the whole test set(which is 1000 in our experiment).
We have several observations:
  1. (1)
    Separate decoder procedure of previous work (Itzhak and Levy, 2022) has seriously underestimated the real knowledge level. As of 7B model, our procedure give an failure rate of 15.0% and separate decoder procedure gives an failure rate of 58.3%. This is exactly what we expect since separate decoder procedure focus on only the embedding layer and ignores the information contained in later layers.
  2. (2)
    In the best estimate (our procedure), we find that Qwen-chat(Bai et al., 2023) failed in nearly 15% Chinese spelling test, and the failure rates seem not to vanish as model size increases, which indicates that the knowledge in token-character relationship of LLMs are clearly limited and incomplete.
The incomplete knowledge in token-character relationship is one of the key reasons why LLM has difficulty in the ”token planning” task. Without further solutions, a token-free model would be neccessary to overcome the difficulty in character-sensitive tasks like poetry writing.
4. Experiment2: Can token-free models do better?
Since the token-based models have been proved to be defective in token-character knowledge, one of the solutions is to use a token-free model. In this section, we will demonstrate how we tailor existing token-based LLMS into token-free models and preserves its pretrained ability to some extent. And we will show how this tailored token-free model could enhance the format accuracy in poetry generation. The detailed token-free model framework is shown in Figure 1.
4.1. Method
Model tailoring. We can certainly pretrain a token-free chinese language model from scratch; on the other hand, we can do better by taking advantage of existing token-based models, tailoring them into token-free models. In this section, we show how this can be done. We can tailor any token-based model into token-free model (in terms of Chinese) by modifying the inputs part and the outputs part of the model.
Inputs. For inputs, we remove all long-tokens from the vocabulary, to ensure that Chinese text will be tokenized into only character-level or byte-level fragments. Specifically, we remove two types of tokens: tokens with more than one Chinese character, and tokens with one Chinese character combined with non-Chinese characters. With those tokens removed, their embeddings in the input layer will never be accessed nor updated.
Note that some keywords in qwen-chat are in English, and those tokens should not been removed. (For example, ”user” and ”assistant” to indicate conversation side are kept in our procedure).
Outputs. For outputs, we add a mask to the output logits (a.k.a language model head), to ensure that the probabilities of long-tokens will always be zero.
Task design. We hope the model could deal with complex task such as story paraphrasing. The final user interface of our model is shown in Figure 2.
User instructions are composed of a format requirement and a content requirement.
Format requirement. Any format type of SHI or CI could be choosen. In the demonstration of our AI Poet (Figure 2), we provide only 12 format types for interface simplicity.
Same as the GPT-4 case in the Introduction section, one example poem of the required format type will be provided in the prompt. To eliminate the possible interference from the content of the example, we replace specific Chinese characters with a [MASK] sign. Thus the imitation task is reformulated as a mask-filling task.
Content requirement. Content instructions are based on one of the following five prompt types:
  1. (1)
    Theme
  2. (2)
    Key words
  3. (3)
    Imagery (A rhetoric used in Chinese classical poetry, which refers to the use of descriptive language that appeals to the senses to create mental pictures or sensations)
  4. (4)
    Language style
  5. (5)
    Main content (A story / A description of scene, etc…)
Following the Qwen-chat prompt format, our prompt for poetry generation is designed as follows.
Prompt $<|im\_start|>$You are a helpful assistant. $<|im\_end|>$ $<|im\_start|>$user Fill in all the $<|extra\_1|>$. {prompt_type}: {prompt} 诗词: {masks}$<|im\_end|>$Äpgjy Response $<|im\_start|>$assistant {fillings}$<|im\_end|>$ $<|endoftext|>$
An example of training datset is provide below. Here the model is required to write a poem in the form of ”Rumengling” and telling a story based on the description.
Prompt $<|im\_start|>$You are a helpful assistant. $<|im\_end|>$ $<|im\_start|>$user Fill in all the $<|extra\_1|>$. 主要内容: 经常记得在溪亭游玩至日暮时分,沉醉其中不知道回家的路。兴致尽头时晚归舟只能迷失,错误地进入了藕花丛深处。争相渡过河流,惊起了一群鸥鹭。 诗词: $<|extra\_1|><|extra\_1|><|extra\_1|>$ $<|extra\_1|><|extra\_1|><|extra\_1|>$ $<|extra\_1|><|extra\_1|><|extra\_1|>$ $<|extra\_1|><|extra\_1|><|extra\_1|>$ $<|extra\_1|><|extra\_1|><|extra\_1|>$ $<|extra\_1|><|extra\_1|>$ $<|extra\_1|><|extra\_1|><|extra\_1|>$ $<|extra\_1|><|extra\_1|><|extra\_1|>$ $<|extra\_1|><|extra\_1|>$ $<|extra\_1|><|extra\_1|>$ $<|extra\_1|><|extra\_1|><|extra\_1|>$ $<|extra\_1|><|extra\_1|><|extra\_1|>$ $<|im\_end|>$Äpgjy Response $<|im\_start|>$assistant 常记溪亭日暮 沉醉不知归路 兴尽晚回舟 误入藕花深处 争渡 争渡 惊起一滩鸥鹭 $<|im\_end|>$ $<|endoftext|>$
4.2. Results
Figure 2. Demonstration of our AI Poet.
Format Type
GPT-4
Jiuge-GPT-2 (token-free, finetuned)
Qwen-chat (token-based, finetuned)
Qwen-chat (token-free, finetuned)
Wuyan Jueju
0.49
1.00
0.79
0.98
Wuyan Lvshi
0.29
1.00
0.94
0.97
Qiyan Jueju
0.88
1.00
0.96
0.98
Qiyan Lvshi
0.81
1.00
0.95
0.97
Rumengling
0.13
0.90
0.85
0.98
Jianzimulanhua
0.81
0.96
0.91
1.00
Qingpingyue
0.13
0.96
0.77
0.95
Dielianhua
0.21
0.91
0.85
1.00
Manjianghong
0.07
0.83
0.78
0.94
Qinyuanchun
0.00
0.55
0.61
0.83
Avg
0.38
0.91
0.84
0.96
Table 1. Format Accuracy of 10 format types, including 4 types of SHI and 6 types of CI.
In the following experiments, we use Qwen-chat-7B as our base model.
Token-free model without further finetune. Surprisingly, without further finetune, the tailored token-free model based on Qwen-chat-7B can already understand humans’ instructions on some daily tasks and generate appropriate answers, though it could not produce poetry with desired format. Below are two examples.
User: 简要介绍一下日心说Äpgjy Agent: 日心说,也称日心地动说,是16th century末期由哥白尼提出的天体学说,它将日、月、水星、金星、木星、土星、天王星、海王星和冥王星等9大天体,及小天体如彗星、小星体等,都绕着以日为圆心的椭圆轨转。
User: Briefly introduce the heliocentric theory.Äpgjy Agent: The heliocentric theory is a model of the solar system that places the Sun at the center, with the planets orbiting around it in椭ical paths. This theory was proposed by Nicolaus Copernicus in the 16th century, but it was not widely accepted until the work of later astronomers such as Johannes Kepler and Galileo Galilei. The heliocentric model is now the most widely accepted model of the solar system, and it is supported by a vast amount of observational evidence.
The first observation is that, without removing non-Chinese tokens, the ability in English conversation is basically left unchanged, though mix-use of Chinese and English occasionally occur.
The second observation is that, ability in Chinese are kept to some extend, even the token representation of Chinese texts are completely different. The model makes some mistakes, while the main idea is correct.
Token-free model finetuned on poetry instruction dataset. We further fintune the tailored model with instruction poetry instruction dataset. The finetuning dataset we use contains 10000 SHI and 10000 CI, formated with the above template. The finetuned model is marked as ”Qwen-chat-7B (token-free, finetuned)”. To verify the effect of token-free method, we also finetune the original token-based Qwen on the exact same dataset as the baseline, marked as ”Qwen-chat-7B (token-based, finetuned)”.
Content Relevance. An example model output comparable to the GPT-4 case in the Introduction section is provided below. Here the model is required to write a poem in the form of ”Rumengling” and with the keyword ”cheerful”. In the example, the model generates a poem translated as ”Talking about the autumn weather, sharing thoughts with others, in high spirits and delight. Ask no more, ask no more. Drinking too much, losing consciousness, and not knowing where I am.”
Response 笑口频开深院,更说秋风天气。 心事向人知,却好兴高采烈。 休觅,休觅,酒到不知醒地。
We can see that the generated poem is relevant in content. In fact, we test on 1000 prompts, with 100 keywords and 10 format styles. Human labelers find no significant difference between the generation quality between the finetuned token-free model and the token-based model finetuned on the same dataset. Our model can also generate poem based on story (story paraphrasing). Examples are provided in the Appendix.
Format Accuracy. Our token-free model perform better in format accuracy at all of the 10 format types, including 4 types of SHI and 6 types of CI. As shown in Table 2, the overall accuracy is 0.96, more than 10% higher compared to the token-based baseline model (0.84). This confirms our hypothesis that token-based model has limited token-character relationship knowledge and that token-free model can generate Chinese poetry with better format accuracy.
Our model also performs better than GPT-4 (0.38) and Jiuge-GPT-2 (0.91). GPT-4 is a general-purpose LLM; Jiuge-GPT-2 is specially designed for poetry generation and is token-free, yet it is not tailored from LLM, and thus could not inherit the great language power from LLM.
5. Conclusion
In this paper, we investigate why the Chinese poem generated by LLMs often violate format requirement. We focus on the token planning problem, and put forward the token-free solution.
We first probe Qwen-chat to show that its knowledge in token-character relationship in Chinese is limited, which is the key reason why they perform bad at poetry format. We formulate the probing procedure as a Supervised-Finetune spelling bee task. Our conclusion is that the knowledge of token-character relationship is not enough for exact token planning.
Then we show that a token-based model can be easily tailored into a token-free model. In regard to input layer and output layer, long-tokens should be removed and only character-level and byte-level tokens should be kept.
We train our token-free model on the poetry generation instructions dataset. Our model is based on Qwen-chat-7B. The tailored model achives an format accuracy of 0.96, much higher than its token-based equivalent (0.84), and also higher than GPT-4 (0.38).